Find the midpoint M of the line segment joining the points S = (-8,2) T and =(0,-4) .
Answer by math_tutor2020(3816)


Add up the x coordinates. Then divide that sum in half.
Add: -8 + 0 = -8
Divide in half: -8/2 = -4
This is the x coordinate of the midpoint.

Repeat the same idea for the y coordinates.
Add: 2+(-4) = -2
Divide in half: -2/2 = -1
This is the y coordinate of the midpoint.


Answer: M = (-4,-1)
You can confirm this answer using GeoGebra or Desmos